# Statistical Demography Meets Ministry of Health: The Case of the Family Planning Estimation Tool

### v1

@@ Abstract
This paper summarizes the main features of the statistical model used in FPET and introduces recent updates related to capturing contraceptive transitions, fitting to survey data that may be error prone, and the use of service statistics data. We use our experience with FPET to discuss lessons learned and open challenges related to the broader field of statistical modeling for monitoring of demographic and global health indicators to help further optimize the application of statistical modeling in practice.
@@ 1 Introduction
The elevation of family planning (FP) on the global stage in 2012 with the launch of FP2020 provided an opportunity to create new approaches for monitoring of FP programs. The need for annual estimates to track progress of the initiative led to the development of standard indicators, approaches, and methodologies. For countries, having annual estimates of FP indicators was a priority to better gauge progress towards their own objectives. After the FP2020 initiative concluded, the FP2030 initiative was started with a continued focus on measurement and setting of country-specific commitments.
@@
Track20, implemented by Avenir Health and funded by the Gates Foundation, is a global family planning project aimed at improving global and country level use of data. The project’s aims include the development and introduction of innovative and user-friendly methodologies, tools, and approaches that build capacity and broaden who can effectively engage with family planning data. Track20 has provided the annual estimates used for FP2020 related measurements and continues to assess progress for FP2030 and the Ouagadougou Partnership (focused on FP in countries in western Africa, see https://www.speakupafrica.org/program/speak-up-africa-website-blurb-for-the-ouagadougou-partnership-op/). This is done through a bottom-up reporting process led by government monitoring and evaluation (M&E) officers, effectively linking country and global FP data.
@@
The Family Planning Estimation Tool (FPET) is used by Track20 in low- and middle-income countries to produce estimates and short-term forecasts of FP indicators, such as modern contraceptive use and unmet need for contraceptives. Estimates are obtained via a Bayesian statistical model that is fitted to country-specific data from surveys and service statistics data. The model is a country-specific implementation of a global model for FP estimation, referred to as the global Family Planning Estimation Model (Alkema et al. [2013], Cahill et al. [2018], Kantorová et al. [2020]). The country-specific implementation was produced to introduce and facilitate in-country usage by the Track20 project. FPET has evolved over the last decade based on usage in the Track20 project and other user inputs and requests. Recent statistical modeling updates include relaxation of strong parametric assumptions made regarding contraceptive use transitions (Susmann and Alkema [2023]), updated use of survey data to improve estimation in the presence of data outliers (Alkema et al. [2024a]), and updated use of service statistics data to better account for uncertainty associated with these data (Mooney et al. [2024a]).
@@
In this paper, we present the current version of the Family Planning Estimation Model and Tool. We first introduce the how FPET came about, how it is used in the Track20 project, and how the modeling approach has evolved. The methods section presents the current model, highlighting how different advances in statistical modeling were incorporated. We use our experience with FPET to discuss lessons learned and open challenges related to the broader field of statistical modeling for monitoring demographic and global health indicators.
@@ 2.1 Why a model is needed to produce FP estimates
Suppose that we are interested in the proportion of married or in-union women aged 15-49 within a population who are using (or whose partner is using) a modern contraceptive method, abbreviated as mCPR. Data on mCPR can be obtained from household surveys in which participants report whether they or their partner are currently using a modern contraceptive method. Such household surveys may be conducted by international organizations or by local organizations or governments. Survey programs and categorizations include the Demographic and Health Survey Program (DHS), Performance Monitoring for Action (PMA), UNICEF Multiple Indicator Cluster Surveys (MICS), national surveys, and other surveys. Micro data from each survey can be used to calculate estimates of mCPR for the population-time period covered by the survey, taking into account the complex sampling design of each survey, as well as a measure of the uncertainty in the mCPR estimate owed to the survey design, referred to as sampling error.
Often survey data alone for some indicator of interest may not be sufficient to inform programs and understand progress. First, data are not necessarily available for all years of interest, e.g., for years in the past, the period after the most recent data point until the current year, and future years. Moreover, data are subject to measurement error, which may be substantial. This issue is discussed in detail in Alkema et al. [2024a]. An example is shown in Figure 1 for Burundi. In Burundi, the most recent national survey, from 2012, suggests a large fluctuation in mCPR which is likely to be due to measurement error. This example illustrates the need to integrate data sources to produce reliable estimates and forecasts.
Figure 1: Data for mCPR among married women for Burundi. Data are shown in colored dots with 95% confidence intervals that reflect sampling errors.
2.2 What FPET does
FPET produces estimates of FP indicators using available data for a population of interest, see Figure 2. In FPET, the FP indicators considered are contraceptive use and unmet need for contraceptives. Methods are categorized into modern versus traditional methods. Modern methods of contraception include female and male sterilization, oral hormonal pills, the intra-uterine device, male and female condoms, injectables, the implant (including Norplant; Wyeth-Ayerst, Collegeville, PA, USA), vaginal barrier methods, standard days method, lactational amenorrhea method, and emergency contraception. Traditional methods of contraception include abstinence, the withdrawal method, the rhythm method, douching, and folk methods. Unmet need for family planning is defined as the percentage of women who want to stop or delay childbearing but who are not currently using any method of contraception to prevent pregnancy. Also included are women who are currently pregnant or postpartum amenorrheic whose pregnancies were mistimed or unwanted.
@@ 2.2 What FPET does
FP indicators are estimated for women aged 15 to 49 years old, considering women’s marital status. More precisely, estimates are produced separately for women who are married or in a union (referred to as married women in this paper), versus women who are not married. Sexual activity is considered in the calculation of the indicators, as summarized in Kantorová et al. [2020].
Data on FP are available from household surveys and routine data collection. As mentioned in the previous section, contraceptive prevalence data obtained from surveys refers to the percentage of women who report themselves or their partners as currently using at least one contraceptive method of any type (modern or traditional). Service statistics data refer to data obtained from routine data collection. These systems may record various FP-related outcomes, including FP commodities distributed to clients or facilities, visits to FP facilities/providers, or family planning users. Track20 has developed a tool that converts service statistics data into a single metric representing volume of services, Estimated Modern Use (EMU), discussed further in a later section.
FPET produces estimates of FP indicators using available survey and service statistics data for a population (e.g., a country or subnational region). An illustration is given in Figure 2 for Kenya.
3 Usage of FPET in the Track20 project
3.1 Overview of Track20’s approach
The Track20 project works to build the capacity of countries to generate and use data to inform their programming, improve impact, and accelerate progress toward their FP goals. One of the pillars of the Track20 approach is the cultivation of a network of M&E Officers dedicated to increasing the quality and use of FP data. These M&E Officers work within or in partnership with Ministries of Health in 34 countries. This approach positions governments to be the driving force of FP data instead of relying on implementing partners to produce needed analyses and findings. Track20 provides ongoing on-demand technical support to these countries. This includes an annual training that brings all M&E Officers together to learn additional analytical skills, engage in south-to-south learning, and to identify opportunities for new analyses and tools based on current government challenges. This direct exchange with multiple countries enables Track20 to develop tools that directly speak to current government priorities. It also provides an opportunity for any new tools to be vetted and tried across many countries, providing valuable feedback that improves Track20’s work.
@@ 3.2 Evolution of FPET
The estimation of FP indicators in FPET has evolved based on user inputs and further model developments. The first version was based on a global modeling approach to estimate and project FP indicators for married women aged 15-49 in all countries in the world using survey data (Alkema et al. [2013]). For in-country usage in the Track20 project, we produced FPET based on a local version of the global model, allowing users to fit the model to data from one country only (New et al. [2017]). This version expanded upon the global model by allowing for the use of service statistics data.
@@
Subsequently, FPET evolved to better capture local contexts. Model updates were introduced to improve predictive performance and to better account for survey data quality issues (Cahill et al. [2018]), and to improve the use of service statistics data (Cahill et al. [2021]). The population considered was extended from married women to all women, using a local implementation of the global model for estimation among unmarried women (Kantorová et al. [2020], Guranich et al. [2021]). In 2024, several updates were introduced to further improve FPET’s utility. The next section introduces the FPET model specification based on these updates.
4 FPET statistical model specification and selected findings
4.1 Overview
@@
We start by introducing notation to formalize the problem of producing estimates and forecasts of FP indicators. We are interested in the proportion of women (by marital status) who either use a modern method of contraception, use a traditional method, have an unmet need for any method, or do not have a demand for FP (no need group). We denote these proportions as $\phi_{j}$ for $j=1,\ldots,4$ referring to modern, trad, unmet any, and no need respectively. Unmet need for modern methods is defined as the sum of traditional users and women with an unmet need for any method. Demand $\eta_{1}$ is defined as the sum of modern users, traditional users, and unmet need for any method: $\eta_{1}=\phi_{1}+\phi_{2}+\phi_{3}$. Demand satisfied is defined as the modern users over demand, $\eta_{2}=\phi_{1}/(\phi_{1}+\phi_{2}+\phi_{3})$. When referring to specific population-periods, additional indices are used. In particular, $\phi_{j,c,m,t}$ refers to the $j$-th proportion for country $c$, marital group $m$, and year $t$.
@@
To produce estimates and forecasts for FP indicators, we made assumptions in the global model about (1) how FP indicators may change over time and vary across different populations, and (2) how available data relate to the indicators. In model descriptions, we distinguish between these two sets of assumptions and refer to the first set as the process model and the second set as the data model assumptions. We introduce the process and data models for modern use and unmet need for modern methods below. The appendix includes the specification of the process and data models for traditional use.
@@
4.2 Process model: How do FP indicators vary over time and differ across populations?
@@
In the process model, we consider how FP indicators change over time. The main goal in FPET is to capture long-term and short-term changes. We do so using a Bayesian hierarchical B-spline transition model with autoregressive country-deviations terms, equivalent to an ARIMA(1,1,0) model with time-varying drift.
Over longer time periods, the main assumptions underlying FPET are based on capturing a contraceptive transition, introduced by Alkema et al. [2013], Cahill et al. [2018] and Susmann and Alkema [2023]. Specifically, over longer time periods, we assume that demand for contraceptives increases from low to high levels, where increases start slowly at low demand, accelerate once growth has been initiated, and then rates of change slow down as demand reaches high levels, which we refer to as an asymptote. Similarly, we assume that demand satisfied with modern methods also follows this type of increase. Illustrations of observed trends globally, that support these assumptions, are given in Figure 3. Based on combining the two transitions, we obtain modern contraceptive use (mCPR) and unmet need for modern methods. Resulting estimates show a typical pattern in unmet need whereby unmet need increases at the start of a transition, when supply is lagging demand, followed by a decrease once supply catches up with demand.
@@
The increase in demand $\eta_{1,c,m,t}$, and demand satisfied with modern methods $\eta_{2,c,m,t}$, from low levels to an asymptote are modeled with flexible transition models (Susmann and Alkema [2023]). In this type of model, the rate of change in an indicator of interest $\eta_{k,c,m,t}$ is parameterized using a systematic component and a smoothing component, where the systematic component captures the expected long-term transition from low to high levels of the indicator while the local smoothing component captures deceleration or acceleration as indicated by the data. Specifically, and dropping indices $k$ and $m$ for notation convenience, a transition model for indicator of interest $\eta_{c,t}$ is specified as follows:
$\Delta\text{logit}(\eta_{c,t})=f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})+\varepsilon_{c,t},$ (1)
where the transition function $f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})$ captures the expected rate of change of the logit-transformed indicator and the stochastic smoothing component $\varepsilon_{c,t}$ captures deviations away from the expected rate of change. The specification for $\eta_{c,t}$ is completed through defining its level in a reference year $t^{*}$:
$\text{logit}(\eta_{c,t^{*}})=\Omega_{c}.$ (2)
The specification of the transition function $f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})$ follows that of Susmann and Alkema [2023], see Figure 4. The transition function is parameterized using B-splines, as follows:
$f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})=\sum_{j=1}^{J}\bm{\beta}_{c,j}B_{j}\left(\eta_{c,t}/\lambda_{c}\right),$ (3)
where $B_{j}(\cdot)$ refers to the $j$th spline basis function, $\bm{\beta}_{c}$ is the vector of splines parameters, and $\lambda_{c}$ is the asymptote. An illustration of this transition function is given in Figure 4.
Figure 4: Illustration of the transition model specification for indicator $\eta_{c,t}$. Left: Rate of change (black) as a function of the $\eta_{c,t}$, parameterized using B-splines. The gray lines illustrate a scaled version of the non-zero spline basis functions. Right: Transition of $\eta_{c,t}$ from low to high levels, based on the rate of change from plot on left. Figure adapted from Susmann and Alkema [2023].
@@
$\varepsilon_{c,t}|\varepsilon_{c,t-1},\rho^{(AR)},\sigma_{\varepsilon}\sim N(\rho^{(AR)}\cdot\varepsilon_{c,t-1},\sigma_{\varepsilon}^{2}),$
with unconditional standard deviation $\text{sd}(\varepsilon_{c,t})=\sigma_{\varepsilon}/\sqrt{1-\rho^{(AR)2}}$. Using an AR(1) smoothing term in the expression for the rate of change implies that the model for $\eta_{c,t}$ itself can be seen as an ARIMA(1,1,0) model on the logit scale with time-varying drift term $f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})$.
Sharing information across populations
Our goal is to estimate and forecast FP indicators from 1970 until 2030 for all countries globally, for women of reproductive age, separately by marital/in-union status. However, a country’s data may not be sufficient to estimate FP for the time period of interest. For example, for countries that are at the start of their transition, country data alone do not allow for estimating its asymptotes. Some countries may have very limited data on FP among unmarried women.
To overcome data limitations, we fit a global model in which information is shared across populations using hierarchical models. We fit the model separately for married versus unmarried women, given that transitions differ between these groups. For each model, we group countries based on geo-political characteristics into clusters and subclusters of countries, following the groupings proposed by Bearak et al. [2020]. In addition, for the modeling of FP among unmarried women, we further extend the groupings to account for differences in cultural norms related to FP and sexual activity among unmarried women, following the approach by Kantorová et al. [2020].
The hierarchical models used, for each marital group, are as follows. Three sets of model parameters are given hierarchical priors:
@@
    asymptotes $\lambda_{k,c}$ for $k=1,2$ controlling the limiting value of an indicator;
@@
    spline coefficients $\bm{\beta}_{k,c,j}$ for $k=1,2$ determining the transition function $f$; and
@@
    $\Omega_{k,c}$ for $k=1,2$, the level of an indicator in a reference year
Specifically, hierarchical priors are assigned to transformations of the model parameters, to constrain parameters to appropriate ranges.
@@
Let $\theta_{\gamma}^{(l)}$ denote the value of one of the transformed parameters for a geographic unit indexed by $\gamma$ at hierarchical level $l$. Level $l$ ranges from $l=1$ referring to the global level, to $l=2$ cluster, $l=3$ subcluster, and $l=4$ country-level. We adopt the following setup:
$\theta_{\gamma}^{(l)}\mid\theta^{(l-1)}_{\gamma^{\prime}},\sigma^{(l)}_{\theta}$ $\sim N(\theta^{(l-1)}_{\gamma^{\prime}},\sigma^{(l)}_{\theta}),\text{ for }l>1,l>1.$

### v2

# Bayesian Statistical Modeling in Action for Estimation and Forecasting in Low- and Middle-income Countries: The Case of the Family Planning Estimation Tool

@@ Abstract
In this paper we summarize the main features of the statistical model used in FPET and introduce recent updates related to capturing contraceptive transitions, fitting to survey data that may be error prone, and the use of service statistics data. We assess model performance through a validation exercise and find that FPET is reasonably well calibrated.
We use our experience with FPET to briefly discuss lessons learned and open challenges related to the broader field of statistical modeling for monitoring of demographic and global health indicators.
@@ 1 Introduction
The elevation of family planning (FP) on the global stage in 2012 with the launch of FP2020 provided an opportunity to create new approaches for monitoring of FP programs ([fp202_2014]). The need for annual estimates to track progress of the initiative led to the development of standard indicators, approaches, and methodologies. For countries, having annual estimates of FP indicators was a priority to better gauge progress towards their own objectives. After the FP2020 initiative concluded, the FP2030 initiative was started with a continued focus on measurement and setting of country-specific commitments ([fp2030]).
@@
Track20, implemented by Avenir Health and funded by the Gates Foundation, is a global family planning project aimed at improving global and country level use of data, see https://www.track20.org/. The project’s aims include the development and introduction of innovative and user-friendly methodologies, tools, and approaches that build capacity and broaden who can effectively engage with family planning data. Track20 has provided the annual estimates used for FP2020 related measurements and continues to assess progress for FP2030 and the Ouagadougou Partnership (focused on FP in countries in western Africa, see https://www.speakupafrica.org/program/speak-up-africa-website-blurb-for-the-ouagadougou-partnership-op/). This is done through a bottom-up reporting process led by government monitoring and evaluation (M&E) officers, effectively linking country and global FP data.
@@
The Family Planning Estimation Tool (FPET) is used by Track20 in low- and middle-income countries to produce estimates and short-term forecasts of FP indicators, such as modern contraceptive use and unmet need for contraceptives. Estimates are obtained via a Bayesian statistical model that is fitted to country-specific data from surveys and service statistics data. The model used in FPET is a country-specific implementation of a global model for FP estimation (alkema2013mcpr, cahill2018mcpr, kantorova_estimating_2020). The country-specific implementation, referred to as a local model, was produced to introduce and facilitate in-country usage by the Track20 project. FPET has evolved over the last decade based on usage in the Track20 project and other user inputs and requests. Recent statistical modeling updates include relaxation of strong parametric assumptions made regarding contraceptive use transitions (susmann_flexible_2025), updated use of survey data to improve estimation in the presence of data outliers (alkema_2024dm), and updated use of service statistics data to better account for uncertainty associated with these data (mooney_emu).
@@
In this paper, we present the current version of the Family Planning Estimation Tool. We first introduce how FPET came about, how it is used in the Track20 project, and how the modeling approach has evolved. The methods section presents the current model set up, highlighting how different advances in statistical modeling were incorporated. We use our experience with FPET to discuss lessons learned and open challenges related to the broader field of statistical modeling for monitoring demographic and global health indicators.
@@ 2 Background
2.1 FP indicators
We are interested in family planning among women of reproductive age in different populations of interest. In this paper, we focus on annual FP outcomes at the national level, from 1970 until 2030, for women based on their marital status. We refer to married women as those women who are married or in a formal union [kantorova_estimating_2020]. Unmarried women are the complement of this group and includes women who are separated, divorced, or widowed.
The FP indicators considered are contraceptive use and unmet need for contraceptives. Contraceptive methods are categorized into modern versus traditional methods. Modern methods of contraception include female and male sterilization, oral hormonal pills, the intra-uterine device, male and female condoms, injectables, the implant (including Norplant; Wyeth-Ayerst, Collegeville, PA, USA), vaginal barrier methods, standard days method, lactational amenorrhea method, and emergency contraception. Traditional methods of contraception include abstinence, the withdrawal method, the rhythm method, douching, and folk methods. Unmet need for family planning among married women is defined as the percentage of women who want to stop or delay childbearing but who are not currently using any method of contraception to prevent pregnancy [bradley_revising_2012]. Also included are women who are currently pregnant or postpartum amenorrheic whose pregnancies were mistimed or unwanted. For unmarried women, sexual activity is considered in the calculation of unmet need, as summarized in kantorova_estimating_2020.
FP indicators are combined to summarize demand for FP and demand satisfied with modern methods. Specifically, demand for FP is defined as the sum of modern users, traditional users, and those women unmet need for any method. Demand satisfied with modern methods, or demand satisfied in short, is defined as the share of women with a demand for FP who use a modern method of contraception.
2.2 Survey data
Survey data on FP are obtained in household surveys in which women answer questions related to marriage, pregnancies, sexual activity, and contraceptive use. For example, contraceptive prevalence data obtained from surveys refers to the percentage of women who report themselves or their partners as currently using at least one contraceptive method of any type (modern or traditional). Household surveys may be conducted by international organizations or by local organizations or governments. Survey programs and categorizations include the Demographic and Health Survey Program (DHS), Performance Monitoring for Action (PMA), UNICEF Multiple Indicator Cluster Surveys (MICS), national surveys, and other surveys. Micro data from each survey can be used to calculate FP indicators for the population-time period covered by the survey, as well as a measure of the uncertainty owed to the survey design, referred to as sampling error.
Survey data compilation is carried out by the United Nations Population Division ([molitoris_world_nodate]) and the Track20 project. The data base contains observations on the compositional vector of FP proportions, i.e., contraceptive use of modern and traditional methods, and unmet need for any method. Demand and demand satisfied can be calculated from these outcomes. Examples of available survey data are shown in Figure 1 for Burundi and Ethiopia.
Figure 1: FP data for married women in Burundi and Ethiopia. Data are shown in colored dots with 95% confidence intervals that reflect sampling errors.
@@ 2.3 Why a model is needed to produce FP estimates
Often survey data alone for some indicator of interest may not be sufficient to inform programs and understand progress. First, data are not necessarily available for all years of interest, e.g., for years in the past, the period after the most recent data point until the current year, and future years. Moreover, data are subject to measurement error, which may be substantial.
These points are illustrated in Figure 1. In Burundi, data are not recent – its most recent survey is from 2016 – and the most recent national survey, from 2012, suggests a large fluctuation in modern contraceptive use (mCPR) which is likely to be due to measurement error. In Ethiopia, data from PMA surveys suggest levels that are lower than those observed in the most recent DHS survey. These examples illustrate the need to integrate data sources to produce reliable estimates and forecasts.
2.4 What FPET does and its usage in the Track20 project
FPET produces estimates of FP indicators using available data for a population of interest, see Figure 2. FPET produces estimates of FP indicators using available surveys for a population (e.g., a country or subnational region). In addition to survey data, service statistics data may also be used to inform the estimates. Service statistics data refer to data obtained from routine data collection. These systems may record various FP-related outcomes, including FP commodities distributed to clients or facilities, visits to FP facilities/providers, or FP users. Track20 developed a tool to calculate an Estimated Modern Use (EMU) indicator from these different types of service statistics data [bietsch_estimated_2025, emutrack20, ss2emu_tool]. EMU estimates have been used to inform mCPR estimates in the absence of recent survey data [cahill2018mcpr, mooney_emu].
@@ 2.4 What FPET does and its usage in the Track20 project
The Track20 project works to build the capacity of countries to generate and use data to inform their programming, improve impact, and accelerate progress toward their FP goals. One of the pillars of the Track20 approach is the cultivation of a network of M&E Officers dedicated to increasing the quality and use of FP data. These M&E Officers work within or in partnership with Ministries of Health in 34 countries. This approach positions governments to be the driving force of FP data instead of relying on implementing partners to produce needed analyses and findings. Track20 provides ongoing on-demand technical support to these countries. This includes an annual training that brings all M&E Officers together to learn additional analytical skills, engage in south-to-south learning (referring to countries in the global south sharing experiences and learning from one another, see for example, https://unsouthsouth.org/about/about-sstc/), and to identify opportunities for new analyses and tools based on current government challenges. This direct exchange with multiple countries enables Track20 to develop tools that directly speak to current government priorities. It also provides an opportunity for any new tools to be vetted and tried across many countries, providing valuable feedback that improves Track20’s work.
@@
The estimation of FP indicators in FPET has evolved based on user inputs and further model developments. The first version was based on a global modeling approach to estimate and project FP indicators for married women aged 15-49 in all countries in the world using survey data (alkema2013mcpr). For in-country usage in the Track20 project, we produced FPET based on a local version of the global model, allowing users to fit the model to data from one country only (new_levels_2017). This version expanded upon the global model by allowing for the use of service statistics data.
@@
Subsequently, FPET evolved to better capture local contexts. Model updates were introduced to improve predictive performance and to better account for survey data quality issues (cahill2018mcpr), and to improve the use of service statistics data (cahill_using_2021). The population considered was extended from married women to all women, using a local implementation of the global model for estimation among unmarried women (kantorova_estimating_2020, guranich_fpemlocal_2021). In 2024, several updates were introduced to further improve FPET’s utility. The next section introduces the FPET model specification based on these updates.
3 Methods
3.1 Overview of workflow
@@
We start by introducing notation to formalize the problem of producing estimates and forecasts of FP indicators. We are interested in the proportion of women of reproductive age in some population who either use a modern method of contraception, use a traditional method, have an unmet need for any method, or do not have a demand for FP (no need group). We denote these proportions as $\phi_{j}$ for $j=1,\ldots,4$ referring to modern use, traditional use, unmet need for any method, and women “with no need“ respectively. Unmet need for modern methods is defined as the sum of traditional users and women with an unmet need for any method. Demand $\eta_{1}$ is defined as the sum of modern users, traditional users, and unmet need for any method: $\eta_{1}=\phi_{1}+\phi_{2}+\phi_{3}$. Demand satisfied is defined as the ratio of modern users over those with a demand, $\eta_{2}=\phi_{1}/(\phi_{1}+\phi_{2}+\phi_{3})$.
For FPET, we are interested in FP indicators by country, for different years, and for women based on their marital status. In terms of notation, when referring to specific population-periods, additional indices are used. In particular, $\phi_{j,c,t,m}$ refers to the $j$-th proportion for country $c$, year $t$, and marital group $m$.
The workflow for estimating FP indicators for all countries and by marital status, and for producing local models for fitting, is summarized in Figure 3. We first fit global models separately by marital status; these models describe the evolution of FP indicators over time in each country and are informed by survey data. Details on the global models are provided in the next section.
The main utility of FPET is to allow countries to produce estimates for a specific population by fitting a model to that population’s data alone. We call these ”local models”: models derived from the global models but adapted so they can be fitted using data from a single population. In FPET, local models are implemented by fixing non-country-specific parameters at the point estimates obtained from the global fits. For local models that estimate FP indicators among all women, service statistics data can also be included. Details on local models are introduced in Section 3.3.
Survey dataGlobal model fittingfor married womenGlobal model fittingfor unmarried womenService statistics dataLocal model fitting(by marital group, orfor both marital groups)Data for 1 country
Figure 3: Workflow for estimating FP indicators by marital status and generating local models for country-level fitting. Global models (green) are fitted by marital status, using survey data. Local models (yellow) use information from global models and are fitted to survey data from one country and, optionally, service statistics data.
3.2 FPET global statistical model specification
3.2.1 Overview
@@
To produce estimates and forecasts for FP indicators for each marital group, we made assumptions in the global model about how available data relate to the indicators and how FP indicators may change over time and vary across different populations. More formally, we defined a likelihood function or data model for available data and a process model that defined the parametrization of model parameters. The diagram in Figure 4 illustrates the set up. We introduce the process and survey data models for modern use and unmet need for modern methods below. The appendix includes the specification of the process and data models for traditional use.
[…]
Figure 4: Overview of global model specification. The process model specifies how FP indicators vary over time and differ across populations. The data model specifies how observed survey data relates to FP indicators. Hierarchical distributions are used for country-specific transition model parameters.
@@
3.2.2 Process model: How do FP indicators vary over time and differ across populations?
@@
In the process model, we consider how FP indicators change over time. The main goal in FPET is to capture long-term and short-term changes. Over longer time periods, the main assumptions underlying FPET are based on capturing a contraceptive transition, introduced by alkema2013mcpr, cahill2018mcpr and susmann_flexible_2025. Specifically, over longer time periods, we assume that demand for contraceptives increases from low to high levels, where increases start slowly at low demand, accelerate once growth has been initiated, and then rates of change slow down as demand reaches high levels, which we refer to as an asymptote. Similarly, we assume that demand satisfied with modern methods also follows this type of increase. Illustrations of observed trends globally, that support these assumptions, are given in Figure 5. Based on combining the two transitions, we obtain modern contraceptive use (mCPR) and unmet need for modern methods. Resulting estimates show a typical pattern in unmet need whereby unmet need increases at the start of a transition, when supply is lagging demand, followed by a decrease once supply catches up with demand.
@@
Transitions in FP indicators are modeled using flexible Bayesian hierarchical B-spline transition models (susmann_flexible_2025). In this type of model, the rate of change in an indicator of interest is parameterized using a systematic component and a smoothing component, where the systematic component captures the expected long-term transition from low to high levels of the indicator while the local smoothing component captures deceleration or acceleration as indicated by the data. We use a Bayesian transition model to capture the increase from low levels to an asymptote in demand and demand satisfied with modern methods.
The remainder of this section introduces the set up of a transition model for one indicator $\eta_{c,t}$. The full specification for the transition models for demand and demand satisfied, for women by marital status, is given in Appendix Section 7.3.
Bayesian B-spline transition model
A transition model for indicator of interest $\eta_{c,t}$ is specified as follows:
$g_{1}(\eta_{c,t})$ $=\begin{cases}\Omega_{c},&t=t^{*},\\ g_{1}(\eta_{c,t-1})+f(\eta_{c,t-1},{\lambda}_{c},{\beta}_{c})+\varepsilon_{c,t},&t>t^{*},\\ g_{1}(\eta_{c,t+1})-f(\eta_{c,t+1},{\lambda}_{c},{\beta}_{c})-\varepsilon_{c,t+1},&t<t^{*},\end{cases}$ (1)
where function $g_{1}()=\Phi^{-1}()$ transforms the indicator to the inverse-probit scale and $\Omega_{c}$ is the level of the indicator on that transformed scale at a fixed reference year $t^{*}$. The transition function $f(\eta_{c,t},{\lambda}_{c},{\beta}_{c})$ captures the expected rate of change of the logit-transformed indicator and the stochastic smoothing component $\varepsilon_{c,t}$ captures deviations away from the expected rate of change.
The specification of the transition function $f(\eta_{c,t},{\lambda}_{c},{\beta}_{c})$ follows that of susmann_flexible_2025, see Figure 6. The transition function is parameterized using B-splines, as follows:
$f(\eta_{c,t},{\lambda}_{c},{\beta}_{c})=\sum_{j=1}^{J}\bm{\beta}_{c,j}B_{j}\left(\eta_{c,t}/\lambda_{c}\right),$ (2)
where $\bm{\beta}_{c}$ is the vector of splines parameters, $B_{j}(\cdot)$ refers to the $j$th spline basis function with normalized domain $[0,1]$, and $\lambda_{c}$ is the asymptote. An illustration of this transition function is given in Figure 6.
Figure 6: Illustration of the transition model specification for indicator $\eta_{c,t}$. Left: Rate of change (black) as a function of the $\eta_{c,t}$, parameterized using B-splines. The gray lines illustrate a scaled version of the non-zero spline basis functions. Right: Transition of $\eta_{c,t}$ from low to high levels, based on the rate of change from plot on left. Figure adapted from susmann_flexible_2025.
@@
$\varepsilon_{c,t}|\varepsilon_{c,t-1},\rho_{\varepsilon},\sigma_{\varepsilon}\sim N(\rho_{\varepsilon}\cdot\varepsilon_{c,t-1},\sigma_{\varepsilon}^{2}),$
with unconditional standard deviation $\text{sd}(\varepsilon_{c,t})=\sigma_{\varepsilon}/\penalty 50\sqrt{1-\rho_{\varepsilon}^{2}}$. Using an AR(1) smoothing term in the expression for the rate of change implies that the model for $\eta_{c,t}$ itself can be seen as an ARIMA(1,1,0) model on the logit scale with time-varying drift term $f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})$.
Sharing information across populations using hierarchical models
Our goal is to estimate and forecast FP indicators from 1970 until 2030 or beyond for all countries globally. However, a country’s data may not be sufficient to estimate FP for the time period of interest. For example, and as illustrated in Figure 1, some countries may have very limited data on FP. For countries that are at the start of their transition, country data alone do not allow for estimating its asymptotes.
To overcome data limitations, we fit a global model in which information is shared across populations using hierarchical models. For each model, we group countries based on geo-political characteristics into clusters and subclusters of countries, following the groupings proposed by bearak_unintended_2020. In addition, for the modeling of FP among unmarried women, we further extend the groupings to account for differences in cultural norms related to FP and sexual activity among unmarried women, following the approach by kantorova_estimating_2020.
In the transition model, three sets of model parameters are given hierarchical priors:
@@
    asymptote $\lambda_{c}$ controlling the limiting value of an indicator;
@@
    spline coefficients $\bm{\beta}_{c,j}$ determining the transition function $f$; and
@@
    $\Omega_{c}$, the level of an indicator in a reference year (on the transformed scale).
For a subset of parameters, hierarchical priors are assigned to transformations of the model parameters, to constrain parameters to appropriate ranges. Full specification of hierarchical models are given in the Appendix. Here we summarize the set up using general notation and give one example.
@@
The general hierarchical model set up is as follows. Let $\theta_{\gamma}^{(l)}$ denote the value of one of the transformed parameters for a geographic unit indexed by $\gamma$ at hierarchical level $l=0,1,\ldots,L$. For national-level estimation, level $l$ starts at the global level with $l=0$ and ends at the country level with $l=L$. Additional levels may include the clusters or subclusters of countries. For each parameter, we adopt the following setup:
$\theta_{\gamma}^{(l)}\mid\theta^{(l-1)}_{\gamma^{\prime}},\sigma^{(l)}_{\theta}$ $\sim N\left(\theta^{(l-1)}_{\gamma^{\prime}},\sigma^{2}_{l,\theta}\right),\text{ for }l>1.$